if 5cosx+5cosy=13 then find max.value of 5sinx+12siny=

Grade:12

3 Answers

siddhant loya
11 years ago

this is an invalid question because of cosx and cosy have max value 1 so 5cosx + 5cosy can be maximum 10 not 13. plzz chk

Ashwin Ashutosh Yardi
11 years ago

The problem is wrong i think.....

5 cos x + 12 cos y = 13   is the correct problem

lokesh soni
11 years ago

this question is wrong since max value of cosx or cosy can be 1 .

and if cox=cosy=1

then max value of 5 cosx+5cosy

=5*1+5*1 =10 

which is less than 13